OpenStudy (anonymous):

(4(y+2)^(1/2)-2y(y+2)^(-1/2))/(y+2) simplify

OpenStudy (helder_edwin):

\[ \large \frac{4\sqrt{y+2}-2y(y+2)^{-1/2}}{y+2} \] ??

OpenStudy (anonymous):

\[\frac{ 4(y+2)^{\frac{ 1 }{ 2 }}-2y(y+2)^{\frac{ -1 }{ 2 }} }{ y+2 }\]

OpenStudy (anonymous):

yeah i guess you can say that

OpenStudy (helder_edwin):

u must know this \[ \large a^{-n} =\frac{1}{a^n}\] then \[ \large ...=\frac{4\sqrt{y+2}-\frac{2y}{\sqrt{y+2}}}{y+2} \]

OpenStudy (helder_edwin):

\[ \large =\frac{\frac{4(y+2)-2y}{\sqrt{y+2}}}{y+2}= \frac{\frac{2y+8}{\sqrt{y+2}}}{y+2}= \frac{2y+8}{(y+2)\sqrt{y+2}} \]

OpenStudy (helder_edwin):

\[ \large =\frac{2y+8}{(y+2)^{3/2}} \]